# SLOctolyzer: Fully automatic analysis toolkit for segmentation and feature extracting in scanning laser ophthalmoscopy images

### v1

@@ 2.3.3 Artery-vein-optic disc detection
The AVOD model used 30 SLO images from two cohorts related to pregnancy (i-Test, 15 images) and multiple sclerosis (FutureMS, 15 images). Given the number of segmentation tasks and dataset size, we similarly opted for fine-tuning a pre-trained U-Net deep learning model, this time using a deep ResNet101 30 backbone, previously trained on ImageNet 28. The dataset was split into training (n=20), validation (n=4) and test (n=6) sets, with equal proportion of cohorts (and thus macula- and disc-centred images) in each set. All images were resized to a common pixel resolution of 768 $\times$ 768 for training.
The model was trained to minimise the Dice Focal loss 24, using the Adam 26 optimiser (with default optimiser hyperparameters) for 300 epochs using a batch size of 4, and a learning rate of 5e-4.
@@ 2.4 Measure module
SLOctolyzer’s second module is a suite of tools for measuring retinal vascular parameters, and are based on the features measured when processing colour fundus images using Automorph 12.
We use the whole image as the region of interest for feature measurement, regardless of the location of the SLO image. Additionally, for disc-centred SLO images, we define two concentric rings centred at the optic disc, zones B and C 21. Zone B takes an annulus from one-half diameter away from the optic disc margin to a full diameter away (0.5D - 1D), and zone C measures from the optic disc margin to two diameters away (0D - 2D). See Fig. 1(B) for a visualisation of these zones.
Across the whole image, we measure global features of vessel complexity, density and calibre of the vasculature, using fractal dimension (Minkowski-Bouligand dimension) 31, vessel density (ratio of vessel pixels to all pixels) and global vessel calibre (ratio of vessel pixels to skeletonised vessel pixels. These measurements are not taken for localised zones B and C. For every region of interest (whole image and zones), we measure tortuosity density 32 and local vessel calibre across individual vessel segments, whose endpoints are defined by arteriovenous crossing and bifurcations. Global and local vessel calibre are very similar measures, but the latter provides a more granular approach to computing calibre from individual vessel segments — see supplementary Fig. S7 and supporting text for more details.
All features are measured across artery, vein and binary vessel maps. Finally, for artery and vein vessel maps, we measure central retinal artery and vein equivalents (CRAE/CRVE) using the Knudston approach 33, as well as the arteriole-to-venule (AVR) ratio. While these metrics are typically only measured for zones B and C in disc-centred images, they are also computed across the whole image and thus for macula-centred images too.
@@ 2.5 SLOctolyzer’s pipeline
Fig. 1 describes the analysis pipeline for SLOctolyzer. During segmentation, the image is resized to a common pixel resolution of 768 $\times$ 768 as the segmentation module was trained on this common pixel resolution. This prevents any segmentation error in larger images due to the central artery/vein light reflex 25, which appears as a bright strip in the centre of a vessel, potentially leading to a larger vessel being perceived as two, smaller and parallel ones. Upon resizing, full segmentation of the vasculature, optic disc and fovea is performed. After thresholding with a value of 0.5 to produce binary vessel masks, post-processing is applied to remove small false positive regions, and disconnected vessels are joined to their neighbouring vessels to remove small gaps. The fovea coordinate is defined as the centroid of the fovea binary mask. If the SLO image is disc-centred, the optic disc binary mask is modelled as an ellipse, and it’s diameter is defined as the average of the major and minor axis lengths.
@@
After segmentation, if the SLO image is macula-centred, measurements on the vessel/artery/vein maps are taken across the whole image. If the SLO image is disc-centred, the measurements are taken across the whole image and for zones B and C, which are defined using the computed optic disc diameter. These measurements will be saved out per file, alongside key metadata either stored in the input file, or inferred during analysis (laterality, location, fovea centre, optic disc diameter and centre). Moreover, segmentation masks will be saved out, as well as helpful visualisations to inspect segmentation quality. If running on a batch of images, a collated file will be saved out row-wise compiling all measurements and metadata from all image files analysed. Finally, a folder of visualisations, superimposing the segmentations onto the SLO image will be saved out per file which will aid segmentation quality checking.
@@
While we do not have an automatic way to assess segmentation quality, an additional feature of SLOctolyzer is its ability to correct manual segmentations. Given a saved out segmentation mask of the arteries, veins and optic disc, the user may correct any erroneous areas using ITK-Snap 22 and save the corrected version in the same folder as the original mask. Upon rerunning the pipeline again, it will use this corrected annotation and re-compute the measurements for that image. See supplementary Fig. S8 and Fig. S9 with supporting text for more information on SLOctolyzer’s setup, interface and output.
@@ 2.6 Statistical analysis
Each segmentation model, after model selection based on the performance against their respective validation set, was evaluated on their respective, internal test set. Additionally, for the AVOD model, we used the RAVIR dataset 10 as an external test to measure the segmentation module’s robustness to severe retinal pathology.
We used standard segmentation metrics such as area under the receiver operating characteristic curve (AUC) and the dice similarity coefficient (at a standard threshold of 0.5) to evaluate segmentation performance. We also used mean absolute error (MAE) to measure performance on downstream, retinal vascular parameters fractal dimension and local vessel calibre (in microns, $\mu$m, where pixel length-scales were available and pixels, px, otherwise). We used MAE to measure the error in predicted fovea pixel coordinate (along both axes) and optic disc area.
@@
We also measured the reproducibility of SLOctolyzer on downstream retinal vascular parameters, using the first 60 participants (120 eyes) of the i-Test cohort (whose data was available at the time of analysis). SLO images were captured simultaneous to macula-centred, OCT volume capture when enhanced depth imaging mode was toggled on and off, creating a repeated, albeit unregistered, pair of SLO images per eye. We measured all retinal vascular parameters previously mentioned across the whole image, and report mean absolute error and Pearson, Spearman and intra-class correlation (ICC(3,1)) coefficients for population-based reproducibility analysis. We also report Bland-Altman plots 34 to assess the distribution of residuals.
Additionally, we report the reproducibility of SLOctolyzer at the eye-level using a measure of individual-level measurement noise $\lambda$ 35. We express the variability of each measured feature within an eye in units of the feature’s overall population variability. $\lambda$ is computed per feature, per eye, as the ratio between the standard deviation of within-eye measurements and the standard deviation of between-eye measurements, across the population. For convenience, $\lambda$ is presented as a percentage, with 0 as the optimal value. Note that the population are all white women with a mean (SD) age of 34.7 (5.2) years, at mean (SD) gestation of 35.6 (3.4) weeks, and thus we do not expect there to be huge variation across the population.
Finally, we measured the execution time of the SLOctolyzer’s segmentation module and the whole pipeline on macula- and disc-centred images of different pixel resolutions. We ran each experiment 100 times and report the mean and standard deviation execution time in seconds. Timed experiments were run on a standard Windows laptop with a four year old i5 CPU with 16GB of RAM, but no GPU.
@@ 3 Results
Feature 0.0130 0.1518 6.2778 4.7778 0.0178 0.2081 0.0133 0.1539 638.1667
Table 3: Segmentation results for each segmentation task (top). Mean absolute error across relevant measurements for each segmentation task (bottom). Average local calibre was used here (in pixels). FD, fractal dimension. Cx is the mean absolute error in the horizontal component of the centroid for the fovea (and similarly for Cy.
Figure 2: Segmentation masks for each segmentation model on two examples randomly selected from their test sets.
@@ 3.1 Evaluation
Segmentation results are presented in Table 3 for all three models for their respective internal test sets. All models had a strong AUC score (all greater than 0.98), suggesting the raw probability maps were well calibrated in being able to score highly the relevant pixels related to each of their segmentation tasks, with the fovea and optic disc detection performing best, which was to be expected since these are easier segmentation tasks than vessel detection. Upon thresholding, both binary vessel and AVOD models scored high dice scores for all-vessel detection (binary vessel model: 0.9097, AVOD model: 0.8570), and the AVOD model scored slightly lower dice scores for artery and vein classification (artery, 0.8376; vein, 0.8525). This was unsurprising, since distinguishing between artery and vein is significantly more challenging than differentiating retinal vasculature from retinal tissue. Fig. 2 shows example segmentation outputs from the internal test sets.
@@
The MAE for fovea coordinate detection was 6.3 and 4.8 pixels for the horizontal and vertical direction (Table 3), thus a MAE of 7.9 pixels diagonally. Images had a pixel resolution of 768 $\times$ 768, covering approximately a 9 mm2 field of view, which makes their conversion factor approximately 11.71 microns-per-pixel in both axes. Thus, this diagonal MAE corresponds to 92.7 microns, which is well within the average size of the foveola of 350 microns 36.
@@
Similarly, an MAE of 638 square pixels for optic disc area corresponds to approximately an MAE of 0.08 mm2, which is not clinically significant given the wide distribution of optic disc sizes 37, the observed change across the myopic spectrum 38 and even differences across different devices 39. MAE for local vessel calibre was around 0.2 pixels, which corresponds to approximately 2.32 microns, and is similarly an insignificant error given population distributions for arterioles and venules using confocal SLO imaging 40.
@@
The results of our segmentation module evaluated against the external test set is shown in Table 4. The RAVIR dataset 10 presents a significant challenge due to retinal pathology heavily featuring in the 23 SLO images. The model performs worse compared with the unseen test data which relate to systemic health (dice coefficient averaged across zones: artery, 0.7228; vein, 0.7598; optic disc, 0.9032), but quantitatively performs well for those images with mild pathology. Given the segmentation models were trained on images with grossly normal retinae related to systemic health changes, it is perhaps no surprise that the performance of the model decreases as the extent of retinal pathology increases. Fig. 3 compares the ground truth labels (middle row) with our ones (bottom row), for SLO images of increasing retinal pathology from left to right. While the major vessels are classified correctly for the majority of cases, as the pathology becomes more severe we see smaller and more tortuous vessels ignored, and segmentations of the major vessels misclassified.
@@
AUC Dice FD Width [px] AUC Dice FD Width [px] AUC Dice Area [px2]
Whole 0.9421 0.7180 0.0444 0.6869 0.9558 0.7470 0.0737 0.9791 0.9720 0.9032 788.5000
Zone B 0.9922 0.7300 0.0379 1.0125 0.9935 0.7726 0.0597 1.1823 - - -
Zone C 0.9643 0.7203 0.0440 0.7854 0.9702 0.7598 0.0543 1.1479 - - -
Table 4: Segmentation and feature measurement mean absolute errors of the artery-vein-optic disc (AVOD) segmentation model against the RAVIR external test set. Average local calibre was used here (in pixels). AUC, area under the receiver operating characteristic curve; FD, fractal dimension (non-dimensional).
@@ 3.2 Reproducibility
Table 5 present correlation values for all retinal vascular parameters for the binary vessel segmentation model (all-vessel), and the artery-vein-optic disc model. All correlations showed strong or excellent agreement among features, with fractal dimension and vessel density scoring the highest. Across the cohort, veins appeared as the poorest performing, but not by any significant margin.
@@
MAE Pearson/Spearman ICC(3,1) MAE Pearson/Spearman ICC(3,1) MAE Pearson/Spearman ICC(3,1)
Fractal dimension 0.0077 0.9326/0.9192 0.965 (0.95, 0.98) 0.0048 0.9485/0.9385 0.9734 (0.96, 0.98) 0.0070 0.8940/0.8860 0.9423 (0.92, 0.96)
Vessel density 0.0013 0.9493/0.9513 0.974 (0.96, 0.98) 0.0022 0.9393/0.9371 0.9686 (0.95, 0.98) 0.0014 0.9370/0.9218 0.9665 (0.95, 0.98)
Global calibre [$\mu$m] 1.5811 0.8771/0.8861 0.9293 (0.9, 0.95) 1.5995 0.8425/0.8331 0.9125 (0.87, 0.94) 1.6063 0.8654/0.8692 0.9276 (0.9, 0.95)
Local calibre [$\mu$m] 1.5049 0.8707/0.8815 0.9259 (0.89, 0.95) 1.4654 0.8489/0.8525 0.9165 (0.88, 0.94) 1.4105 0.8695/0.8711 0.9299 (0.9, 0.95)
Tortuosity density 0.0324 0.7431/0.7460 0.8526 (0.79, 0.9) 0.0147 0.7962/0.7977 0.886 (0.84, 0.92) 0.0241 0.7039/0.6825 0.8262 (0.75, 0.88)
AVR/CRAE/CRVE [$\mu$m] 0.0283 0.7743/0.7931 0.8728 (0.82, 0.91) 6.2834 0.8456/0.8554 0.9151 (0.88, 0.94) 10.1025 0.7940/0.8009 0.8852 (0.83, 0.92)
@@
Table 5: Population-based reproducibility performance of SLOctolyzer, reporting mean absolute error (MAE), Pearson, Spearman and intra-class correlations (ICC(3,1)). All Pearson and Spearman correlations were statistically significant with P-values $P<0.0001$. For readability, AVR/CRAE/CRVE were combined into one row — AVR is dimensionless, while CRAE and CRVE are in microns and only report for arteries and veins, respectively.
@@
Fig. 4 present Bland-Altman plots of the residuals between paired SLO images. We found that SLOctolyzer was highly reproducible, with the distribution of residuals according to most metrics centred around 0, showing no apparent trend. For fractal dimension, there was a mean difference (MD) of -0.0007 with limits of agreement (LoA) of [-0.0178, 0.0164]; for vessel density there was a MD of -0.0003 and LoA of [-0.0052, 0.0046]; for global vessel calibre there was a MD of -0.3154 $\mu$m and LoA of [-4.9804, 4.3496]; for local vessel calibre there was a MD of -0.3208 $\mu$m and LoA of [-4.5777, 3.9361]; for tortuosity density there was a MD of -0.0013 and LoA of [-0.0624, 0.0650]; for CRAE there was a MD of 0.9901 $\mu$m and LoA of [-17.4345, 15.4543]; for CRVE there was a MD of -1.8782 $\mu$m and LoA of [-30.3251, 26.5687]; finally, for AVR there was a MD of 0.0011 and LoA of [-0.0738, 0.0760].
@@
Figure 4: Bland-Altman analysis of the residuals from assessing the reproducibility of the segmentation models for all features outputted by SLOctolyzer. Residuals for arteries (red), veins (blue) and all-vessel (orange) are plotted together.
@@
Fig. 5 reports the eye-level reproducibility of each of the features measured by SLOctolyzer. For fractal dimension, vessel density and global/local calibre, the upper, or third quartile is below 25% across all vessel types. Thus, the range of measurement error for these features in the majority of the paired, but unregistered, eyes were below 25% of the overall populations’ variability. Given the homogeneous nature of the population’s demographics, this is very reasonable. Moreover, vessel density has the smallest interquartile range among features (mean all-vessels, $\lambda=11.7\%$; arteries, $\lambda=12.1\%$; veins, $\lambda=13.3\%$), suggesting this feature, closely followed by fractal dimension, had the least measurement noise. Tortuosity density and AVR/CRAE/CRVE appear the most sensitive features, with veins as the most variable relative to arteries and all-vessels.
@@
The eyes with the worst reproducibility for vessel calibre and tortuosity density are shown to the right of the box plots, with dashed lines linking the major outlier scatter-points to the corresponding eye. For all calibre metrics, the worst metric came from the same eye. This was due to an arteriovenous crossing near the edge of the image, where a major artery and vessel overlapped, and SLOctolyzer classified artery and vein differently (red arrows). This led to a large absolute difference of 87.5 microns in CRVE. For tortuosity density, the vessel maps do not look qualitatively different, but the absolute difference was 0.096. We believe that this is due to tortuosity density’s sensitivity to arteriovenous crossings which disconnect the respective artery and vein vessel maps. Star-shaped scatter-points represent second to largest outliers for vessel calibre and tortuosity density and are shown in supplementary Fig. S10, and are similar due to arteriovenous crossings or vessel disconnectedness.
@@
Figure 5: Reproducibility at the eye-level of each of the features outputted by SLOctolyzer. Box-plot distributions are shown for each feature, with all-vessels in orange, arteries in red and veins in blue. Note that the AVR/CRAE/CRVE features are presented together as CRVE/CRAE are measured across arteries (red) and veins (blue) individually, while AVR is a combination of the two, thus represented in the all-vessel class (orange). A horizontal, dashed line at the $\lambda$=25mark is shown as a visual aid. We also show the eyes (with segmentations overlaid) which had the greatest reproducibility measurement error for large-vessel calibre and vein tortuosity density (circular scatter-points). Eyes corresponding to the star-shaped scatter-points are shown in supplementary Fig. S10.

### v2

@@ 2.3.3 Artery-vein-optic disc detection
Given the number of segmentation tasks and dataset size, we similarly opted for fine-tuning a pre-trained UNet deep learning model, this time using a deep ResNet101 32 backbone, previously trained on ImageNet 30. Details on how to reproduce the model architecture are shown in supplementary code listing LABEL:suppcode:fovea_code and surrounding text.
The AVOD model was trained on a total of 30 SLO images. The dataset was split into training (n=20), validation (n=4) and test (n=6) sets, with equal proportion of cohorts (and thus macula- and disc-centred images) in each set. All images were resized to a common image resolution of 768 $\times$ 768 pixels for training.
The model was trained to minimise the Dice Focal loss 26, using the Adam 28 optimiser (with default optimiser hyperparameters) for 300 epochs using a batch size of 4, and a learning rate of 5e-4.
@@ 2.4 Measure module
Figure 2: (A) Disc-centred SLO image with segmentations and regions of interest overlaid. (B) Region of interest masks for the whole image (green), zone C (blue) and zone B (red) with the names of vessel metric, colour coordinated with arrows indicating which regions they are measured in. (C) Flowchart of which vessel metrics are computed for which segmentation map.
SLOctolyzer’s second module is a suite of tools for measuring retinal vascular parameters and are based on the features measured when processing colour fundus images using Automorph 13. Fig. 2 shows the different regions of interest defined and the measurements taken per region and per vessel map.
We define only one region of interest for macula-centred images and three regions of interest for disc-centred images to compute feature measurements on. The region of interest common across macula- and disc-centred images is the whole image (Fig. 2(B), green). Relevant to only disc-centred images are the second and third regions of interest, which define two concentric rings centred at the optic disc called zones B and C 22. Zone B takes an annulus from one-half diameter (D) away from the optic disc margin to a full diameter away (0.5D – 1D) (Fig. 2(B), red), and zone C measures from the optic disc margin to two diameters away (0D – 2D) (Fig. 2(B), blue).
Fig. 2(B) shows which metrics are measured across which regions of interest. Across the whole image, we measure global features of density, complexity and calibre, computed using vessel density (the ratio of vessel pixels to all pixels), fractal dimension (Minkowski-Bouligand dimension) 33 and global vessel calibre (ratio of vessel pixels to skeletonised vessel pixels) — these measurements are not taken for localised zones B and C. For every region of interest (whole image and zones B and C), we measure local vessel calibre and tortuosity density 34 across individual vessel segments whose endpoints are defined by arteriovenous crossing and bifurcations. Global and local vessel calibre are very similar metrics, but the latter provides a more granular approach to computing calibre from individual vessel segments — see supplementary Fig. S10 and supporting text.
Fig. 2(C) shows a simple flowchart of which measurements are taken across which vessel map. All of these aforementioned features are measured across artery, vein and all-vessel maps (Fig. 2(C), green). For artery and vein vessel maps only, we measure central retinal artery and vein equivalents (CRAE and CRVE) using the Knudston approach 35, respectively, as well as the arteriole-to-venule (AVR) ratio (Fig. 2(C), red-blue-purple). Note that although CRAE/CRVE/AVR are typically only measured for zones B and C in disc-centred images, they are also computed across the whole image and for macula-centred images too.
@@ 2.5 SLOctolyzer’s pipeline
Fig. 1 describes the analysis pipeline for SLOctolyzer. Please see supplementary Fig. S11 and Fig. S12 with supporting text for more information on SLOctolyzer’s setup, interface and output.
2.5.1 Input
SLOctolyzer supports macula-centred and disc-centred SLO images inputted as regular image files (.jpeg, .png, etc.), as well as .vol file formats (from Heidelberg Engineering imaging devices). In the former, the end-user can optionally input information on spatial sampling, laterality (right/left) and location (macula-/disc-centred). If the laterality and location of the SLO image is not inputted, these are inferred based on the detected locations of the fovea and optic disc after segmentation, and measurements are outputted in pixel space. For inputting .vol files, these store the image data and metadata necessary for converting measurements from pixel space to physical space — other file formats such as .e2e and .dcm are currently not supported, but we are working on this presently. A configuration file is also provided for the end-user to specify input/output directories, whether the pipeline skips over files which have unexpected errors during processing (helpful for large-scale batch processing), and the option to ignore saving out segmentation maps (to reduce memory consumption during processing).
2.5.2 Segmentation & measurement
@@
During segmentation, the image is first resized to a common image resolution of 768 $\times$ 768 pixels as the segmentation module was trained on this image resolution, thus making the pipeline invariant to different spatial sampling length-scales (assuming a fixed field of view). Moreover, this resizing prevents any segmentation error in images with larger image resolution due to the central artery/vein light reflex 27 which appears as a bright strip in the centre of a vessel, potentially leading to a large vessel being perceived as two, smaller and parallel vessels. Upon resizing, full segmentation of the vasculature, optic disc and fovea is performed and the probability segmentation maps are then resized to the image’s original image resolution. After thresholding with a value of 0.5 to produce binary masks, post-processing is applied to remove small false positive regions, and disconnected vessels are joined to their neighbouring vessels to remove small gaps.
The foveal pit is defined as the centroid of the fovea binary mask. If the SLO image is disc-centred, the optic disc binary mask is modelled as an ellipse, and it’s centre and diameter are the ellipse’s centre and the average of its major and minor axis lengths, respectively. If the SLO image is macula-centred, measurements on the all-vessel/artery/vein maps are only taken across the whole image. If the SLO image is disc-centred, measurements are taken across the whole image and for zones B and C, which are defined using the optic disc diameter.
2.5.3 Output
@@
Measurements will be saved out per file, alongside a process log and key metadata stored located in the file’s metadata or inferred during analysis (laterality, location, fovea centre, optic disc diameter and centre). Optionally, segmentation masks will be saved out, as well as helpful visualisations to inspect segmentation quality. If running on a batch of images, a collated .xlsx file will be saved out row-wise compiling all measurements and metadata from all image files analysed. Finally, a folder of visualisations, superimposing the segmentations onto the SLO image will be saved out per file which will aid segmentation quality checking.
@@
While we do not have an automatic way to assess segmentation quality, an additional feature of SLOctolyzer is its ability to correct manual segmentations. Given a saved-out segmentation mask of the arteries, veins and optic disc, the end-user may correct any erroneous areas using ITK-Snap 23 and save the corrected version in the same folder as the original mask. Upon rerunning the pipeline again, it will use this corrected annotation and re-compute the measurements for that image.
@@ 2.6 Statistical analysis
Each segmentation model, after model selection based on the performance against their respective validation set, was evaluated on their respective, internal test set. Additionally, for the AVOD model, we used the RAVIR dataset 11 as an external test to measure the segmentation module’s robustness to severe retinal pathology.
We used standard segmentation metrics such as area under the receiver operating characteristic curve (AUC) and the dice similarity coefficient (at a standard threshold of 0.5) to evaluate segmentation performance. We also used mean absolute error (MAE) to measure performance on downstream, retinal vascular parameters fractal dimension and local vessel calibre (in microns [$\mu$m], where transversal spatial sampling length-scales were available and pixels [px] otherwise). We used MAE to measure the error in predicted fovea pixel coordinate (along both axes) and optic disc area.
@@
We also measured the reproducibility of SLOctolyzer on downstream retinal vascular parameters, using the first 60 participants (120 eyes) of the i-Test cohort (whose data was available at the time of analysis). SLO images were captured simultaneous to macula-centred, OCT volume capture when enhanced depth imaging mode was toggled on and off, creating a repeated, albeit unregistered, pair of SLO images per eye. We measured all retinal vascular parameters previously mentioned across the whole image, and report mean absolute error and Pearson (P), Spearman (S) and intra-class correlation (ICC(3,1)) coefficients for population-based reproducibility analysis. We also report Bland-Altman plots 36 to assess the distribution of residuals.
Additionally, we report the reproducibility of SLOctolyzer at the eye-level using a measure of individual-level measurement noise $\lambda$ 37. We express the variability of each measured feature within an eye in units of the feature’s overall population variability. $\lambda$ is computed per feature, per eye, as the ratio between the standard deviation of within-eye measurements and the standard deviation of between-eye measurements, across the population. For convenience, $\lambda$ is presented as a percentage, with 0 as the optimal value. Note that the population are all white women with a mean (SD) age of 34.7 (5.2) years, at mean (SD) gestation of 35.6 (3.4) weeks, and thus we do not expect there to be huge variation across the population.
[…]
@@ 3 Results
Feature 0.01 0.15 6.28 4.78 0.02 0.21 0.01 0.15 638.17
Table 3: Segmentation results for each segmentation task (top). Mean absolute error across relevant measurements for each segmentation task (bottom). Calibre, local vessel calibre; FD, fractal dimension; Cx/Cy, mean absolute error in the horizontal/vertical component of the centroid for the fovea; px, pixels.
Figure 3: Predicted segmentation masks from each model for two examples randomly selected from their internal test sets.
@@ 3.1 Evaluation
Segmentation results are presented in Table 3 for all three models for their respective internal test sets. All models had a strong AUC score (all greater than 0.98), suggesting the raw probability maps were well calibrated in being able to score highly the relevant pixels related to each of their segmentation tasks, with the fovea and optic disc detection performing best, which was to be expected because these are easier segmentation tasks than vessel detection. Upon thresholding, both binary vessel and AVOD models scored high dice scores for all-vessel detection (binary vessel model: 0.91, AVOD model: 0.86), and the AVOD model scored slightly lower dice scores for artery and vein classification (artery, 0.84; vein, 0.85). This was unsurprising, because distinguishing between artery and vein is significantly more challenging than differentiating retinal vasculature from retinal tissue. Fig. 3 shows example segmentation outputs from the internal test sets.
@@
The MAE for fovea coordinate detection was 6.3 and 4.8 pixels for the horizontal and vertical axes (see Table 3), thus reporting an MAE of 7.9 pixels diagonally. Images had an image resolution of 768 $\times$ 768 pixels, covering approximately a 9 mm2 region of interest, which makes their transversal spatial length-scale approximately 11.71 microns-per-pixel in both axes. Thus, this diagonal MAE corresponds to approximately 92.7 microns, which is well within the estimated 350-micron width of the foveola centralis 24.
@@
Similarly, an MAE of 638 square pixels for optic disc area corresponds to approximately an MAE of 0.08 mm2, which is not clinically significant given the wide distribution of optic disc sizes 38, the observed change across the myopic spectrum 39 and even differences across different devices 40. MAE for local vessel calibre was around 0.2 pixels, which corresponds to approximately 2.32 microns, and is similarly an insignificant error given population distributions for arterioles and venules using confocal SLO imaging 41.
@@
The results of our segmentation module evaluated against the external test set is shown in Table 4. The RAVIR dataset 11 presents a significant challenge due to retinal pathology heavily featuring in the 23 SLO images. The model performs worse compared with the unseen, internal test data which relate to systemic health (dice coefficient averaged across zones: artery, 0.72; vein, 0.76; optic disc, 0.90), but quantitatively performs well for those images with mild pathology. Given the segmentation models were trained on images with grossly normal retinae related to systemic disease, it is perhaps no surprise that the performance of the model decreases as the extent of retinal pathology increases. Fig. 4 qualitatively compares the ground truth labels (middle row) with SLOctolyzer’s predictions (bottom row), for SLO images of increasing retinal pathology from left-to-right. While the major vessels are classified correctly for the majority of cases, as the pathology becomes more severe we see smaller and more tortuous vessels ignored, and segmentations of the major vessels misclassified.
@@
AUC Dice FD Calibre [px] AUC Dice FD Calibre [px] AUC Dice Area [px2]
Whole Image 0.94 0.72 0.04 0.69 0.96 0.75 0.07 0.98 0.97 0.90 788.50
Zone B 0.99 0.73 0.04 1.01 0.99 0.77 0.06 1.18 - - -
Zone C 0.96 0.72 0.04 0.79 0.97 0.76 0.05 1.15 - - -
Table 4: Segmentation and feature measurement mean absolute errors of the artery-vein-optic disc (AVOD) segmentation model against the RAVIR external test set. Calibre, local vessel calibre; AUC, area under the receiver operating characteristic curve; FD, fractal dimension; px, pixels.
@@ 3.2 Reproducibility
Table 5 presents correlation values for all retinal vascular parameters for the binary vessel segmentation model (all-vessel), and the artery-vein-optic disc model. All correlations showed strong or excellent agreement among features, with fractal dimension and vessel density scoring the highest. Across the cohort, veins appeared as the poorest performing, but not by any significant margin.
@@
MAE P/S ICC(3,1) MAE P/S ICC(3,1) MAE P/S ICC(3,1)
Fractal dimension 0.008 0.93/0.92 0.97 (0.95, 0.98) 0.005 0.95/0.94 0.97 (0.96, 0.98) 0.007 0.89/0.89 0.94 (0.92, 0.96)
Vessel density 0.001 0.95/0.95 0.97 (0.96, 0.98) 0.002 0.94/0.94 0.97 (0.95, 0.98) 0.001 0.94/0.92 0.97 (0.95, 0.98)
Global calibre [$\mu$m] 1.58 0.88/0.89 0.93 (0.90, 0.95) 1.60 0.84/0.83 0.91 (0.87, 0.94) 1.61 0.87/0.87 0.93 (0.90, 0.95)
Local calibre [$\mu$m] 1.50 0.87/0.88 0.93 (0.89, 0.95) 1.47 0.85/0.85 0.92 (0.88, 0.94) 1.41 0.87/0.87 0.93 (0.90, 0.95)
Tortuosity density 0.03 0.74/0.75 0.85 (0.79, 0.9) 0.01 0.80/0.80 0.89 (0.84, 0.92) 0.02 0.70/0.68 0.83 (0.75, 0.88)
AVR/CRAE/CRVE [$\mu$m] 0.03 0.77/0.79 0.87 (0.82, 0.91) 6.28 0.85/0.86 0.92 (0.88, 0.94) 10.10 0.79/0.80 0.89 (0.83, 0.92)
@@
Table 5: Population-based reproducibility performance of SLOctolyzer, reporting mean absolute error (MAE), Pearson (P), Spearman (S) and intra-class correlations (ICC(3,1)). All Pearson and Spearman correlations were statistically significant with P-values $P<0.0001$. For readability, AVR/CRAE/CRVE were combined into one row — AVR is dimensionless and specified in the all-vessel column, while CRAE and CRVE are in microns and only report for the artery and vein maps, respectively.
@@
Fig. 5 shows Bland-Altman plots of the residuals between paired, repeated SLO images. We found that SLOctolyzer was highly reproducible, with the distribution of residuals according to most metrics centred around 0, showing no apparent trend. For fractal dimension, there was a mean difference (MD) of -0.001 with limits of agreement (LoA) of [-0.019, 0.017]; for vessel density there was a MD of -0.0003 and LoA of [-0.0052, 0.0046]; for global vessel calibre there was a MD of -0.32 $\mu$m and LoA of [-4.98, 4.34]; for local vessel calibre there was a MD of -0.32 $\mu$m and LoA of [-4.57, 3.93]; for tortuosity density there was a MD of 0.001 and LoA of [-0.062, 0.064]; for CRAE there was a MD of -0.99 $\mu$m and LoA of [-17.43, 15.45]; for CRVE there was a MD of -1.88 $\mu$m and LoA of [-30.32, 26.56]; finally, for AVR there was a MD of 0.001 and LoA of [-0.073, 0.075].
@@
Figure 5: Bland-Altman plots of residual distributions from assessing the reproducibility of SLOctolyzer’s segmentation module across all retinal vessel metrics. Scatter-plot residuals for arteries (red), veins (blue) and all-vessel (orange) are overlaid together.
@@
Fig. 6 reports the eye-level reproducibility of each of the features measured by SLOctolyzer. For fractal dimension, vessel density and global/local calibre, the upper, or third quartile is below 25% across all vessel types. Thus, the range of measurement error for these features in the majority of the paired, but unregistered, eyes were below 25% of the overall populations’ variability. Given the homogeneous nature of the population’s demographics, this is very reasonable. Moreover, vessel density has the smallest interquartile range among features (mean all-vessels, $\lambda=11.7\%$; arteries, $\lambda=12.1\%$; veins, $\lambda=13.3\%$), suggesting this feature, closely followed by fractal dimension, had the least measurement noise. Tortuosity density and AVR/CRAE/CRVE appear as the most sensitive features, with veins as the most variable relative to arteries and all-vessels.
@@
The eyes with the worst reproducibility for vessel calibre and tortuosity density are shown to the right of the box plots, with dashed lines linking the major outlier scatter-points to the corresponding eye’s SLO image. For all calibre metrics, the worst metric came from the same eye. This was due to an arteriovenous crossing near the edge of the image, where a major artery and vessel overlapped, and SLOctolyzer classified artery and vein differently (red arrows). This led to a large absolute difference of 87.5 microns in CRVE. For tortuosity density, the vessel maps do not look qualitatively different, but the absolute difference was 0.096. We believe that this is due to tortuosity density’s sensitivity to arteriovenous crossings which disconnect the respective artery and vein vessel maps. Star-shaped scatter-points represent second to largest outliers for vessel calibre and tortuosity density and are shown in supplementary Fig. S13, and are similarly due to arteriovenous crossings or vessel disconnectedness.
@@
Figure 6: Reproducibility at the eye-level of each of the features outputted by SLOctolyzer. Box-plot distributions are shown for each feature, with all-vessels in orange, arteries in red and veins in blue. Note that the AVR/CRAE/CRVE features are presented together as CRVE/CRAE are measured across arteries (red) and veins (blue) separately, while AVR is a combination of the two, thus represented in the all-vessel class (orange). A horizontal, dashed line at the $\lambda$=25mark is shown as a visual aid. We also show the eyes (with segmentations overlaid) which had the greatest eye-level reproducibility measurement error for large-vessel calibre and vein tortuosity density (circular scatter-points). Eyes corresponding to the star-shaped scatter-points are shown in supplementary Fig. S13.